OpenStudy (anonymous):

For the equation xy + 4x + 7y = −4, calculate the derivative dy/dx at the point (1, −1).

OpenStudy (bahrom7893):

x*(dy/dx)+y + 4 + 7(dy/dx) = -4 Plug in x = 1, y = -1 (dy/dx) - 1 + 4 + 7(dy/dx) = -4 8(dy/dx) = -7 dy/dx = -7/8

OpenStudy (anonymous):

Bahroom do you care to help me all I have left is 2 more?

OpenStudy (anonymous):

Sorry for interupting

OpenStudy (anonymous):

its -3/8

OpenStudy (bahrom7893):

then i probably added or subtracted wrong. just check the arithmetic... oh yea the derivative of a constant is 0, so: x*(dy/dx)+y + 4 + 7(dy/dx) = 0 (dy/dx) - 1 + 4 + 7(dy/dx) = 0 8(dy/dx) = -3 dy/dx = -3/8

OpenStudy (anonymous):

thanks man!